OpenStudy (anonymous):

grocery sold 17 products. Total is $8.65. How many at 0.65 and how many at 0.45 ?

OpenStudy (anonymous):

x*0,65 + (17-x)*0,45 = 8,65 x*0,65 + 17*0,45 -x*0,45 = 8,65 x*0,2 = 8,65 - 7,65 x*0,2 = 1 x = 1/0,2 = 5 So number of products sold at 0,65 is 5 and number of products sold at 0,45 is 12.

OpenStudy (anonymous):

^ except the other was at 0.65, that's incorrect.

OpenStudy (anonymous):

x*0,45 + (17-x)*0,65 = 8,65 x*0,45 + 17*0,65 -x*0,65 = 8,65 11,05 - 8,65 = 0,2*x 2,4 = 0,2*x x = 12 (number of 0,45) 5 (number of 0,65) Same answer my friend...